OpenStudy (anonymous):

Solve for the real value of x: 1/x+x=4

OpenStudy (anonymous):

x= 3.73, & x= 0.27

OpenStudy (anonymous):

how did you get that answer i'm very confused and have a test monday

OpenStudy (anonymous):

I'll help. So the equation is \[(1/x) +x=4\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Alrighty then.. in order to solve it you have to simplify it in order to get rid of the denominator. So inorder to do this you need to multiply x to both sides. \[x((1/x)+x)=(4)x\]

OpenStudy (anonymous):

Once you do that you should get \[1+x^2=4x\]

OpenStudy (anonymous):

Sorry I can explain another time. I have some studying to do. And thanks mods for the warning, how about next time talking to me first. Much Love

OpenStudy (anonymous):

Then you move everything to the left side of the equation and you get \[x^2-4x+1=0\] Since you can't factor it you have to use the quadratic formula which is \[x=(-b \pm \sqrt{b^2-4ac})/2a\] and you insert the numbers into the formula with 1=a, -4=b and 1=c. and you get \[x=(-(-4)\pm \sqrt{(-4)^2-4(1)(1)})/2(1)\]

OpenStudy (anonymous):

from there you simplify. \[x=\frac{ 4\pm \sqrt{16-4} }{ 2 }\] \[x=\frac{ 4\pm \sqrt{12} }{ 2}\] \[x=\frac{ 4\pm2\sqrt{3}}{ 2 }\] \[x=2\pm \sqrt{3}\]

OpenStudy (anonymous):

in this instance is the quadratic the only equation you can use ?

OpenStudy (anonymous):

Yes. Since you can't factor it out by hand or by multiplying two numbers to get the c value and adding to get the b values.

OpenStudy (anonymous):

thank you so much i'm starting to understand now !

OpenStudy (anonymous):

Your welcome. I apologize if it's not all that clear to you.